OpenStudy (anonymous):

Josie makes fruit punch by mixing fruit juice and lemonade in the ratio 2:3, She needs to make 30 litres of punch for a party. How much of each ingedient does she need? Fruit juice = ? Lemonade= ?

OpenStudy (ganpat):

Fruit Juice = 30 * 2/5 Lemonade = 30 * 3/5...

OpenStudy (anonymous):

why is 2/5??can u explain

OpenStudy (anonymous):

is that how many litres of fruit juice and lemonade?

OpenStudy (ganpat):

if u take 2/ 3 raito.. than it turns out to be 2in proportion of 5 for juice and 3 in proportion of 5 for lemonade..

OpenStudy (anonymous):

okay i got it thx

OpenStudy (ganpat):

so juice = 30 * 2/5 = 12 liters Lemonade = 30 * 3/5 = 18 liters.. @moha_10 : from the above, when u add, it comes to be 30 liters.. so the proportion needs to be added as 2 + 3....

OpenStudy (anonymous):

ok thanks:)

OpenStudy (ganpat):

W/C

OpenStudy (anonymous):

@Ganpat thx